OpenStudy (anonymous):

Find all the Real Solutions for the following equations. \[x^3 - 3x = y\] \[y^3 - 3y=z\] \[z^3 - 3z = x\]

OpenStudy (anonymous):

(z^2 - 3) (x^2 - 3) (y^2 - 3) = 1

OpenStudy (anonymous):

x (x^2 - 3) = y x^2 - 3 = y/x y (y^2 - 3) = z y^2 - 3 = z/y z (z^2 - 3) = x z^2 - 3 = x/z (z^2 - 3) (x^2 - 3) = y/z (z^2 - 3) (x^2 - 3) = 1/ (y^2 - 3) (z^2 - 3) (x^2 - 3) (y^2 - 3) = 1